# Skagit Fisheries Enhancement Group   UMass Amherst

### v1

@@ 4 ALDI++: Improving DAOD
We next propose two novel enhancements to the Align and Distill approach, resulting in a new method ALDI++. We show in Sec. 6.2 that these enhancements lead to state-of-the-art results, and ablate each component in Sec. 6.3.
1. Robust burn-in. First we propose a new “burn-in” strategy for pretraining a teacher model $\theta_{tch}$ on source-only data $X_{src}$. A key challenge in student-teacher methods is improving target-domain pseudo-label quality. We point out that psuedo-label quality in the early stages of self-training is largely determined by the out-of-distribution (OOD) generalization capabilities of the initial teacher model $\theta^{init}_{tch}$, and thus propose a training strategy aimed at improving OOD generalization during burn-in. We add strong data augmentations including random resizing, color jitter, and random erasing, and keep an EMA copy of the model during burn-in, two strategies that have previously been shown to improve OOD generalization and robustness [1, 2, 18]. We are the first to utilize these strategies for DAOD burn-in.
2. Multi-task soft distillation. Most prior work utilizes confidence thresholding and non-maximum suppression to generate “hard” pseudo-labels from teacher predictions $\hat{p}_{tgt}$ (see Tab. 1). However in object detection this strategy is sensitive to the confidence threshold chosen, leading to both false positive and false negative errors that harm self-training [31]. We take inspiration from the knowledge distillation literature and propose instead using “soft” distillation losses—i.e. using teacher prediction scores as targets without thresholding—allowing us to eliminate the confidence threshold hyperparameter.
We distill each task of Faster R-CNN—Region Proposal Network localization ($rpn$) and objectness ($obj$), and Region-of-Interest Heads localization ($roih$) and classification ($cls$)—independently. At each stage, the teacher provides distillation targets for the same set of input proposals used by the student—i.e. anchors $A$ in the first stage, and student region proposals $p^{rpn}_{tgt}$ in the second stage: $p^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{stu}(A,x^{t}_{tgt})$ (1) $\hat{p}^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{tch}(A,x^{\hat{t}}_{tgt})$ (2) $p^{roih,cls}_{tgt}=\theta^{roih,cls}_{stu}(p^{rpn}_{tgt},x^{t}_{tgt})$ (3) $\hat{p}^{roih,cls}_{tgt}=\theta^{roih,cls}_{tch}(p^{rpn}_{tgt},x^{\hat{t}}_{tgt})$ (4)
@@
Where $L_{rpn}$ and $L_{roih}$ are the smooth L1 loss and $L_{obj}$ and $L_{cls}$ are the cross-entropy loss, and $\lambda_{0\dots 3}=1$ by default. See Fig. 2 for a visual depiction. We include more implementation details in the supplemental material.
One prior DAOD work, PT [8], has also used soft distillation losses, however we note two shortcomings our method addresses: (1) PT requires a custom “Probabilistic R-CNN” architecture for distillation, while our approach is general and can work with any two-stage detector, and (2) PT uses $\hat{p}^{cls}$ as an indirect proxy for distilling $p^{obj}$, while our approach is able to distill each task directly.
@@ 5 The CFC-DAOD Dataset
Next we introduce our dataset contribution, CFC-DAOD, addressing P3: (a) Lack of diverse benchmarks leading to overestimation of methods’ generality.
@@
CFC. The Caltech Fish Counting Dataset (CFC) [30] is a domain generalization benchmark sourced from fisheries monitoring, where sonar video is used to detect and count migrating salmon. The detection task consists of a single class (“fish”) and domain shift is caused by real-world environmental differences between camera deployments. We identify this application as an opportunity to study the generality of DAOD methods due to its stark differences with existing DAOD benchmarks—specifically, sonar imagery is grayscale, has low signal-to-noise ratios, and foreground objects are difficult to distinguish from the background—however CFC focuses on generalization rather than adaptation and does not include the data needed for DAOD.
CFC-DAOD We introduce an extension to CFC, deemed CFC-DAOD, to enable the study of DAOD in this application domain. The task is to adapt from a source location—“Kenai”, i.e. the default training set from CFC—to a difficult target location, “Channel”. We collected an additional 168k bounding box annotations in 29k frames sampled from 150 new videos captured over two days from 3 different sensors on the “Channel” river (see Fig. 3). For consistency, we closely followed the video sampling protocol used to collect the original CFC dataset as described by the authors (see [30]). Our addition to CFC is crucial for DAOD as it adds an unsupervised training set for domain adaptation methods and a supervised training set to train oracle methods. We keep the original supervised Kenai training set from CFC (132k annotations in 70k images) and the original Channel test set (42k annotations in 13k images). We note this is substantially larger than existing DAOD benchmarks (CS contains 32k instances in 3.5k images, and Sim10k contains 58k instances in 10k images). See the supplemental material for more dataset statistics. We make the dataset public.
@@ 6 Experiments
In this section we propose an updated benchmarking protocol for DAOD (Sec. 6.1) that allows us to fairly analyze the performance of ALDI++ compared to prior work (Sec. 6.2) and conduct extensive ablation studies (Sec. 6.3).
Datasets. We perform experiments on Cityscapes $\rightarrow$ Foggy Cityscapes, Sim10k $\rightarrow$ Cityscapes, and CFC Kenai $\rightarrow$ Channel. In addition to being consistent with prior work, these datasets represent three common adaptation scenarios capturing a range of real-world challenges: weather adaptation, Sim2Real, and environmental adaptation, respectively. We note that there have been inconsistencies in prior work in terms of which ground truth labels for Cityscapes are used. We use the Detectron2 version.
Metrics. For all experiments we report the PascalVOC metric of mean Average Precision with IoU $\geq 0.5$ (“AP50”) [15]. This is consistent with prior work on Cityscapes, Foggy Cityscapes, Sim10k, and CFC.
6.1 A New Benchmarking Protocol for DAOD
@@
Revisiting source-only and oracle models. Here we address P1: Improperly constructed source-only and oracle models, leading to overestimation of performance gains. The goal of DAOD is to develop adaptation techniques that use unlabeled target-domain data to improve target-domain performance. Thus, in order to properly isolate adaptation-specific techniques, any technique that does not need target-domain data to run should also be used by source-only and oracle models. In our case, this means that source-only and oracle models should also utilize the same strong augmentations and EMA updates as DAOD methods. In Fig. 4 we illustrate the resulting source-only and oracle models, and show that including these components significantly improves both source-only and oracle model performance (+7.2 and +2.6 AP50 on Foggy Cityscapes, respectively). This has significant implications for DAOD research: because source-only and oracle models have not been constructed with equivalent components, performance gains stemming from better generalization have until now been misattributed to DAOD. With properly constructed source-only and oracle models, the gains from DAOD are much more modest (see Fig. 5).
Modernizing architectures. We next address P3: (b) Outdated model architectures leading to overestimation of methods’ generality. Prior art in DAOD has used older backbones (e.g. VGG-16) in order to compare to previously-published results. To investigate whether conclusions drawn from these results generalize to modern experimental settings, our experiments utilize a modern detection framework [54] with default settings including multi-scale input transforms and COCO pre-training. We use a ResNet-50 backbone [23] with Feature Pyramid Network [35] and ViTDet [33]. We provide more details in the supplementary. Source-only and oracle models also receive these upgrades.
@@ 6.2 Fair Comparison and State-of-the-Art Results
We compare ALDI++ with reimplementations of five state-of-the-art DAOD methods on top of our framework: UMT [12], SADA [10], PT [8], MIC [26], and AT [34]; see Tab. 1 for the ALDI settings used to reproduce them. We use the fair benchmarking protocol proposed in Sec. 6.1. Results are shown in Fig. 5. All methods (including ALDI++) use the same settings for all benchmarks.
ALDI++ is state-of-the-art on CS normal-$\rightarrow$ FCS, Sim10k normal-$\rightarrow$ CS, and CFC Kenai normal-$\rightarrow$ Channel. ALDI++ outperforms the previous state-of-the-art by +3.5 AP50 on CS $\rightarrow$ FCS, +5.7 AP50 on Sim10k $\rightarrow$ CS (where ours is the only method to outperform a fair source-only model), and +2.0 AP50 on CFC Kenai $\rightarrow$ Channel. Further, we achieve near-oracle level performance on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel (0.4 and 0.9 AP50 away, respectively), while other methods close less than half the gap between source-only and oracle models.
Modern architectures and fair source-only models create a paradigm shift for DAOD benchmarking. Our benchmarking protocol provides a dramatic reset for DAOD performance bounds. A fair and modern source-only model—trained without ever seeing any target domain data—achieves a higher target AP50 than all previously-published DAOD methods (see Fig. 1). Similarly, we see a 57% increase in oracle performance compared to prior work.
@@
Figure 5: ALDI++ outperforms all prior work on CS normal-$\rightarrow$ FCS, Sim10k normal-$\rightarrow$ CS, and CFC Kenai normal-$\rightarrow$ Channel. We provide a fair comparison of ALDI++ with five existing state-of-the-art approaches by upgrading them to use the ALDI framework. We see that some prior methods continue to provide benefit on top of a modern architecture but others lag behind modern source-only models. Our method ALDI++ outperforms prior work on all datasets studied by a significant margin: +3.5 AP50 on CS normal-$\rightarrow$ FCS, +5.7 AP50 on Sim10k normal-$\rightarrow$ CS, and +2.0 AP50 on CFC Kenai normal-$\rightarrow$ Channel. Notably, ALDI++ is the only method to outperform a source-only model on Sim10k normal-$\rightarrow$ CS.
Relative performance of all methods decreases compared to fair source-only and oracle models. Re-implementing SOTA methods in ALDI improves absolute performance of all methods; however performance decreases compared to source-only models. There are several instances where modernized DAOD methods are actually worse than a fair source-only model. Notably, a source-only model outperforms upgraded versions all previously-published work on Sim10 $\rightarrow$ CS. We also see that no state-of-the-art methods outperform a fair oracle on any dataset, in contrast to claims made by prior work [34, 8, 5].
@@
The ranking of methods differs across datasets and architectures. MIC and AT are consistently the top-performing methods across all datasets. UMT exhibits variable performance due to the differences in the difficulty of image generation across datasets (see supplemental for examples). SADA underperforms other methods on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel, but closes this gap on the more difficult Sim10k $\rightarrow$ CS. These differences highlight the importance of benchmarking in a modern context, as we see that previously-published methods are not always complementary to general advancements in object detection. These results also demonstrate the utility of CFC-DAOD as another point of comparison for DAOD methods.
ALDI++ is compatible with new detector variants. We upgrade ALDI++ to use VitDet [33]. Since VitDet is a two-stage architecture based on Faster R-CNN, this requires no modifications to our multi-task distillation loss. We show that ALDI++ continue to show improvements over an upgraded VitDet source-only model (see Fig. 1 for CS $\rightarrow$ FCS and supplemental for other datasets). We see there is a larger gap between the ViT ALDI++ and the ViT oracle, indicating the potential for future work to improve performance.
@@ 6.3 Ablation Studies
In this section we ablate the performance of each component of ALDI on CS $\rightarrow$ FCS. For each ablation, unless otherwise specified we begin with the settings shown in Tab. 2, ablating one column at a time. Our default burn-in uses the same augmentations as the Fixed strategy from Tab. 1 but uses early stopping for model selection, and all self-training runs are initialized with the same burned-in checkpoint for fair comparison. See Tab. 1 for additional definitions.
Table 2: Starting point for all ablation experiments.
Method $\theta_{stu},\theta_{tch}$ (Burn-in) $T_{src}$ $T_{tgt}$ $\frac{B_{tgt}}{B}$ Postprocess $L_{distill}$ $L_{align}$
Ablation start Weak augs, early stopping F, M F, M, J, C 0.5 Thresh, NMS Hard
@@
$\theta_{stu}$, $\theta_{tch}$ Network initialization (burn-in). In Fig. 6a we analyze the effects of our proposed burn-in strategy (see Sec. 4). We measure performance in terms of target-domain AP50 as well as convergence time, defined as the training time at which the model first exceeds 95% of its final target-domain performance. We compare our approach with: (1) No dataset-specific burn-in, i.e. starting with COCO weights, and (2) The approach used by past work—using a fixed burn-in duration, e.g. 10k iterations. We find that our method results in significant improvements in both training speed and accuracy, leading to upwards of 10% improvements in AP50 and reducing training time by a factor of 10 compared to training without burn-in.
[…]
@@
$T_{src}$ Source augmentations. In Fig. 6b we ablate the set of source-domain data augmentations. We compare using weak augmentations (random flipping and multi-scale training), strong augmentations (color jitter and random erasing), and a combination of weak and strong, noting that prior works differ in this regard but do not typically report the settings used (see Tab. 1). We find that using strong source augmentations on the entire source-domain training batch outperforms weak augmentations and a combination of both.
@@
$T_{tgt}$ Target augmentations. In LABEL:tab:target_aug we investigate the use of different augmentations for target-domain inputs to the student model. (We note that weak augmentations are always used for target-domain inputs to the teacher in accordance with prior work). We see that stronger augmentations consistently improve performance, with best performance coming from the recently-proposed MIC augmentation [26].
@@
Table 3: (a) Effects of target-domain augmentation on self-training. Augmentations applied to student inputs ($T_{tgt}$ in Fig. 2). Stronger augmentations improve performance considerably. (b) Effects of distillation objectives on self-training. We compare hard targets—used by most prior art, which thresholds teacher predictions to create pseudo-labels—with our proposed soft targets. Soft targets can improve overall performance. Results are the mean and 95% confidence intervals over 3 runs. (c) Feature alignment has diminishing returns. Alignment objectives $L_{align}$ without self-training provides up to 2.6 AP50 of benefit (left column), but diminishes to 0.2 AP50 additional gain when used alongside self-training (right column).
[…]
Source-only model 51.9

### v3

@@ 4 ALDI++: Improving DAOD
We next propose a set of simple but effective enhancements to the Align and Distill approach. We call the resulting method ALDI++. We show in Section 6.2 that these enhancements lead to state-of-the-art results, and ablate each component in Section 6.3.
1. Robust burn-in. A key challenge in student-teacher methods is improving target-domain pseudo-label quality. We point out that pseudo-label quality in the early stages of self-training is largely determined by the out-of-distribution (OOD) generalization capabilities of the initial teacher model $\theta^{init}_{tch}$, and thus propose a pre-training (“burn-in”) strategy aimed at improving OOD generalization before self-training.
We add strong data augmentations including random resizing, color jitter, and cutout (DeVries & Taylor, 2017; Chen et al., 2020), and keep an EMA copy of the model during burn-in, two strategies that have previously been shown to improve OOD generalization (Morales-Brotons et al., 2024; Arpit et al., 2022; Gao et al., 2022), i.e. we pre-train a model $\theta$ with the loss from Eq. 1, where $t\sim T_{\text{src}}$ and ${L}_{\text{sup}}$ are still the standard Faster R-CNN losses. Each iteration we update an EMA copy of the model,
$\theta^{\text{EMA}}=\alpha\theta^{\text{EMA}}+(1-\alpha)\theta$ (6)
with $\alpha\in[0,1]$. After pre-training, we initialize $\theta_{\text{stu}}=\theta_{\text{tch}}=\theta^{\text{EMA}}$. We are the first to utilize these strategies for DAOD burn-in, and we show in Section 6.3 that this pre-training strategy leads to faster convergence time and better results.
2. Multi-task soft distillation. Most prior work utilizes confidence thresholding and non-maximum suppression to generate “hard” pseudo-labels from teacher predictions $\hat{p}_{tgt}$. However in object detection this strategy is sensitive to the confidence threshold, leading to false positive and false negative errors that harm self-training (Kay et al., 2023; RoyChowdhury et al., 2019). Inspired by the knowledge distillation literature we propose instead using “soft” distillation losses—i.e. using teacher prediction scores as targets without thresholding—allowing us to eliminate the confidence threshold hyperparameter.
We describe here our approach for two-stage (Faster R-CNN-based) object detection. Distillation implementation details for YOLO and DETR architectures can be found in Appendix B. We distill each task of Faster R-CNN—Region Proposal Network localization ($rpn$) and objectness ($obj$), and Region-of-Interest Heads localization ($roih$) and classification ($cls$)—independently. At each stage, the teacher provides distillation targets for the same set of input proposals used by the student—i.e. anchors $A$ in the first stage, and student region proposals $p^{rpn}_{tgt}$ in the second stage:
$p^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{stu}(A,x^{t}_{tgt})$ (7)
$\hat{p}^{rpn,obj}_{tgt}=\theta^{rpn,obj}_{tch}(A,x^{\hat{t}}_{tgt})$ (8)
$p^{roih,cls}_{tgt}=\theta^{roih,cls}_{stu}(p^{rpn}_{tgt},x^{t}_{tgt})$ (9)
$\hat{p}^{roih,cls}_{tgt}=\theta^{roih,cls}_{tch}(p^{rpn}_{tgt},x^{\hat{t}}_{tgt})$ (10)
@@
Where $L_{rpn}$ and $L_{roih}$ are the smooth L1 loss and $L_{obj}$ and $L_{cls}$ are the cross-entropy loss, and $\lambda_{0\dots 3}=1$ by default. See Fig. 2 for a visual depiction, and the appendix for implementation details.
One prior DAOD work, PT (Chen et al., 2022), has also used soft distillation losses. Our method addresses two shortcomings: (1) PT requires a custom “Probabilistic R-CNN” architecture for distillation, while our approach is general and can work with any two-stage detector, and (2) PT uses $\hat{p}^{cls}$ as an indirect proxy for distilling $p^{obj}$, while our approach distills each task directly.
3. Revisiting DAOD training recipes. We also re-examine common design choices in DAOD in order to establish strong baseline settings for ALDI++. In particular, we find that two simple changes consistently improve domain adaptation results: (1) Using strong regularization on both target and source data during self-training, and (2) Training with equal amounts of source and target supervision in each minibatch (i.e. $B_{src}=B_{tgt}$). We also opt to disable all feature alignment in ALDI++ to stabilize training and find that the effects on accuracy are minimal (see Section 6.3).
@@ 5 The CFC-DAOD Dataset
Next we introduce our dataset contribution, CFC-DAOD, as a step toward addressing P3: (a) Lack of diverse benchmarks leading to overestimation of methods’ generality.
CFC. The Caltech Fish Counting Dataset (CFC) (Kay et al., 2022) is a domain generalization benchmark sourced from fisheries monitoring, where sonar video is used to detect and count migrating salmon. The detection task consists of a single class (“fish”) and domain shift is caused by real-world environmental differences between camera deployments. We identify this application as an opportunity to study the generality of DAOD methods due to its stark differences with existing DAOD benchmarks—specifically, sonar imagery is grayscale, has low signal-to-noise ratios, and foreground objects are difficult to distinguish from the background—however CFC focuses on generalization rather than adaptation and does not include the data needed for DAOD.
CFC-DAOD We introduce an extension to CFC, deemed CFC-DAOD, to enable the study of DAOD in this application domain. The task is to adapt from a source location—“Kenai”, i.e. the default training set from CFC—to a difficult target location, “Channel”. We collected an additional 168k bounding box annotations in 29k frames sampled from 150 new videos captured over two days from 3 different sensors on the “Channel” river (see Fig. 3). For consistency, we closely followed the video sampling protocol used to collect the original CFC dataset as described by the authors (see (Kay et al., 2022)). Our addition to CFC is crucial for DAOD as it adds an unsupervised training set for domain adaptation methods and a supervised training set to train oracle methods. We keep the original supervised Kenai training set from CFC (132k annotations in 70k images) and the original Channel test set (42k annotations in 13k images). We note this is substantially larger than existing DAOD benchmarks (CS contains 32k instances in 3.5k images, and Sim10k contains 58k instances in 10k images). See Appendix D for more dataset statistics and Appendix G for qualitative visualizations. We make the dataset public.
@@ 6 Experiments
In this section we propose an updated benchmarking protocol for DAOD (Section 6.1) that allows us to fairly analyze the performance of ALDI++ compared to prior work (Section 6.2) and conduct extensive ablation studies (Section 6.3).
6.1 Benchmarking Protocol
Datasets. We perform experiments on Cityscapes $\rightarrow$ Foggy Cityscapes, Sim10k $\rightarrow$ Cityscapes, and CFC Kenai $\rightarrow$ Channel. In addition to being consistent with prior work, these datasets represent three common adaptation scenarios capturing a range of real-world challenges: weather adaptation, Sim2Real, and environmental adaptation, respectively. We note that there have been inconsistencies in prior work in terms of which ground truth labels for Cityscapes are used. We use the Detectron2 version, which includes three intensity levels of fog $\{0.005,0.01,0.02\}$.
Metrics. For all experiments we report the PascalVOC metric of mean Average Precision with IoU $\geq 0.5$ (“AP50”) (Everingham et al., 2010). This is consistent with prior work on Cityscapes, Foggy Cityscapes, Sim10k, and CFC.
Revisiting source-only and oracle models. Here we address P1: Improperly constructed source-only and oracle models, leading to overestimation of performance gains. The goal of DAOD is to develop adaptation techniques that use unlabeled target-domain data to improve target-domain performance. Thus, in order to properly isolate adaptation-specific techniques, any technique that does not need target-domain data to run should also be used by source-only and oracle models. This means that source-only and oracle models should also utilize the same strong augmentations and EMA updates as DAOD methods.
In Fig. 4 we illustrate the resulting source-only and oracle models, and show that including these components significantly improves both source-only and oracle model performance (+7.2 and +2.6 AP50 on Foggy Cityscapes, respectively). This has significant implications for DAOD research: because source-only and oracle models have not been constructed with equivalent components, performance gains stemming from better generalization have until now been misattributed to DAOD. With properly constructed source-only and oracle models, the gains from DAOD are much more modest (see Fig. 5). Note that for clarity we compare all methods in Fig. 1 and Fig. 5 against a single source-only and oracle model, however it would be more appropriate to compare each method to its own bespoke source-only and oracle models that use the same training components; see Section A.7 for the full comparison.
Fixed training settings. Prior work has used inconsistent backbones and image sizes, making head-to-head comparisons less fair. Using ALDI we instead compare using the same training settings, offering a fair comparison. As a starting point we utilize reasonably modern settings likely to be used by a practitioner: the Cityscapes defaults in the Detectron2 codebase. All methods in our comparisons, including source-only and oracle models, utilize Faster R-CNN architectures with ResNet-50 (Ren et al., 2015) backbones with FPN (Lin et al., 2017), COCO (Lin et al., 2014) pre-training, and an image size of 1024px on the shortest side. See Appendix C for more details.
@@ 6.2 Fair Comparison and State-of-the-Art Results
Figure 5: Fair comparison of ALDI++ with existing state-of-the-art approaches using the ALDI framework and modern training recipes. Some prior methods show consistent benefit but others lag behind fair source-only models. Our method ALDI++ outperforms prior work on all datasets studied by a significant margin: +3.5 AP50 on CS $\rightarrow$ FCS, +5.7 AP50 on Sim10k $\rightarrow$ CS, and +0.6 AP50 on CFC Kenai $\rightarrow$ Channel. Notably, ALDI++ is the only method to outperform a source-only model on Sim10k $\rightarrow$ CS.
We compare ALDI++ with reimplementations of five state-of-the-art DAOD methods on top of our framework: UMT (Deng et al., 2021), SADA (Chen et al., 2021), PT (Chen et al., 2022), MIC (Hoyer et al., 2023), and AT (Li et al., 2022b); see Appendix C for the ALDI settings used to reproduce them. We use the fair benchmarking protocol proposed in Section 6.1. Results are shown in Fig. 5. All methods (including ALDI++) use the same settings for all benchmarks.
Comparison with state-of-the-art. ALDI++ outperforms all prior work and sets a new state-of-the-art on all benchmarks studied, outperforming the next-best methods by +3.5 AP50 on CS $\rightarrow$ FCS, +5.7 AP50 on Sim10k $\rightarrow$ CS (where ours is the only method to outperform a fair source-only model), and +0.6 AP50 on CFC Kenai $\rightarrow$ Channel. ALDI++ achieves near-oracle level performance on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel (0.4 and 0.9 AP50 away, respectively), while other methods close less than half the gap between source-only and oracle models.
@@
Comparison across datasets. We compare all methods on CS $\rightarrow$ FCS, Sim10k $\rightarrow$ CS, and CFC Kenai $\rightarrow$ Channel, in Fig. 5. We find the ranking of methods differs across datasets. ALDI++, MIC and AT are consistently the top-performing methods across all datasets. MIC performs noticeably better on CFC Kenai $\rightarrow$ Channel than other prior work, nearly matching the performance of ALDI++. UMT exhibits variable performance due to the differences in the difficulty of image generation across datasets (see Appendix C for examples). SADA underperforms other methods on CS $\rightarrow$ FCS and CFC Kenai $\rightarrow$ Channel, but closes this gap on the more difficult Sim10k $\rightarrow$ CS. These results demonstrate the utility of CFC-DAOD as another point of comparison for DAOD methods; we see that method performance on synthetic benchmarks like CS $\rightarrow$ FCS is not necessarily indicative of performance on real-world domain shifts.
Comparison with fair source-only and oracle models. Re-implementing methods in ALDI improves absolute performance of most methods due to upgraded training settings; however performance decreases dramatically compared to source-only and oracle models. There are several instances where modernized DAOD methods are actually worse than a fair source-only model. Notably, a source-only model outperforms upgraded versions all previously-published work on Sim10 $\rightarrow$ CS. We also see that no state-of-the-art methods outperform a fair oracle on any dataset, in contrast to claims made by prior work (Li et al., 2022b; Chen et al., 2022; Cao et al., 2023).
Figure 6: Comparison across backbones and pre-training data. Pre-training strategy does not significantly change the relative strength of methods compared to each other nor compared to source-only and oracle models. However, all models perform worse compared to source-only and oracle models when using VitDet backbones, with only ALDI++ and MIC outperforming a source-only model.
Comparison across backbones and pre-training data. We compare all methods on CS $\rightarrow$ FCS using ImageNet vs. COCO pre-training, as well as Resnet-50 FPN backbones vs. VitDet-B backbones, in Fig. 6. See Appendix F for other datasets. We see that while COCO pre-training improves absolute performance of all methods, their ranking does not change significantly compared to ImageNet pre-training, nor does their performance in relation to source-only baselines. Interestingly, we find that ALDI++ outperforms an oracle on CS $\rightarrow$ FCS when using ImageNet pre-training, however we do not observe this trend on the other datasets (see Appendix F). We hypothesize that this may be due to noise in the Foggy Cityscapes target-train labels due to their programmatic generation, and that COCO pre-training helps prevent the oracle from overfitting to these erroneous boxes. We show that only ALDI++ and MIC continue to show improvements over an upgraded VitDet source-only model (see Section A.2 for other datasets), with ALDI++ performing slightly better than MIC (+0.4 AP50). We see there is a larger gap between the ViT ALDI++ and the ViT oracle compared to ResNet backbones, indicating the potential for future work to improve performance. Across all experiments in Fig. 6 we see that the source-only–oracle gap shrinks as the underlying model improves due to either stronger pre-training or backbone upgrades, indicating that DAOD may offer diminishing returns with stronger models.
@@ 6.3 Ablation Studies
In this section we ablate the performance of each component of ALDI on CS $\rightarrow$ FCS.
Base settings. For each ablation, unless otherwise specified we begin with the following training settings. We initialize $\theta_{stu},\theta_{tch}$ with COCO pre-training followed by a burn-in phase on $X_{\text{src}}$ with weak augmentations and early stopping based on validation performance. $T_{\text{src}}$ includes random horizontal flip and random scaling. $T_{\text{tgt}}$ includes random horizontal flip, random scaling, color jitter, and cutout. The $B_{\text{src}}:B_{\text{tgt}}$ batch ratio is 1:1. $L_{\text{distill}}$ is hard pseudo-labeling with a confidence threshold of 0.8, and $L_{\text{align}}$ is disabled. Note these base settings are not necessarily those of ALDI++ but rather the most commonly chosen values in prior work for each component. Additional training settings are reported in Appendix C.
@@
$\theta_{stu}$, $\theta_{tch}$ Network initialization (burn-in). In Fig. 7a we analyze the effects of our proposed burn-in strategy (see Section 4). We measure performance in terms of target-domain AP50 as well as convergence time, defined as the training time at which the model first exceeds 95% of its final target-domain performance. We compare our approach with: (1) No dataset-specific burn-in, i.e. starting with COCO weights, and (2) The approach used by past work—using a fixed burn-in duration, e.g. 10k iterations. We find that our method results in significant improvements in both training speed and accuracy, leading to upwards of 10% improvements in AP50 and reducing training time by a factor of 10 compared to training without burn-in.
Figure 7: Ablation studies. (a) Our proposed burn-in strategy (Section 4) improves AP50FCS by +4.7 and reduces training time by 10x compared to no burn-in. (b) Strong source-data augmentations during self-training lead to better performance. (c) An equal ratio of source and target data during self-training leads to best performance.
@@
$T_{src}$ Source augmentations. In Fig. 7b we ablate the set of source-domain data augmentations. We compare using weak augmentations (random flipping and random scaling), strong augmentations (color jitter and cutout), and a combination of weak and strong, noting that prior works differ in this regard but do not typically report the settings used. We find that using strong source augmentations on the entire source-domain training batch outperforms weak augmentations and a combination of both.
@@
$T_{tgt}$ Target augmentations. In Table 1a we investigate the use of different augmentations for target-domain inputs to the student model. (We note that weak augmentations are always used for target-domain inputs to the teacher in accordance with prior work). We see that stronger augmentations consistently improve performance, with best performance coming from the recently-proposed MIC augmentation (Hoyer et al., 2023).
@@
Table 1: Ablation studies. (a) Effects of target-domain augmentation on self-training. Augmentations applied to student inputs ($T_{tgt}$ in Fig. 2). Stronger augmentations improve performance considerably. (b) Effects of distillation objectives on self-training. We compare hard targets—used by most prior art, which thresholds teacher predictions to create pseudo-labels—with our proposed soft targets. Soft targets can improve overall performance. Results are the mean and standard deviation over 3 runs. (c) Feature alignment has diminishing returns. Compared to a source-only baseline AP50 of 59.1, feature alignment objectives $L_{align}$ without self-training provides up to 2.6 AP50 of benefit (first row), but diminishes to 0.2 AP50 additional gain when used alongside self-training (last row).
$T_{tgt}$ AP50FCS